OpenStudy (anonymous):

how do you do this integral: ∫(t^2)sin⁡(t^3+t^4)√(2t+9t^4+16t^6 )dt from 0 to 5?

OpenStudy (anonymous):

That problem is insane

OpenStudy (anonymous):

yah i know

OpenStudy (anonymous):

can't use calculator either

OpenStudy (anonymous):

i typed it into wolframalpha.com it solved the problem but it didnt give the result

OpenStudy (anonymous):

I get - 12.01 but after the 01 the number goes even further than that

OpenStudy (anonymous):

i dont know if thats the right answer

OpenStudy (anonymous):

unless there is another way around it. original problem is fine the line integran \[\int\limits_{}^{} xsin(y+z)ds

OpenStudy (anonymous):

oops lemme do that again

OpenStudy (anonymous):

find the line integral: \[\int\limits_{}^{} xsin(y+z)ds\]

OpenStudy (anonymous):

where x = t^2 y= t^3 z=t^4 \[0\le t \le5\]

OpenStudy (anonymous):

the answer should be around 15 too

OpenStudy (anonymous):

hmm im not sure

OpenStudy (anonymous):

thats what the professor told us.

OpenStudy (anonymous):

wow not much of a professor lol I thought professor knew everything

OpenStudy (anonymous):

well he told us he wanted us to solve it. And that the answer was around 15 to kinda help us i guess. But no clue where to even begin solving that integral

OpenStudy (anonymous):

wow i still dont know where to start

OpenStudy (anonymous):

yeah same

OpenStudy (anonymous):

i think you can use integration by parts here. http://en.wikipedia.org/wiki/Integration_by_parts

OpenStudy (anonymous):

umm.. not sure how i would approach it with that

OpenStudy (anonymous):

well, I don't remember the procedure now. I just know that these type of integrations can be solved using integration by parts. I think the wikipedia page has the description of how to do it. Sorry I can't be of more help.

OpenStudy (anonymous):

ok, thanks anyways.